Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate are as follows: Standard Hours Standard Rate per Hour Standard Cost 27 minutes $5.40 $2.43 During August, 9,275 hours of direct labor time were needed to make 19,300 units of the Jogging Mate. The direct labor cost totaled $49,158 for the month. Required: 1. What is the standard labor-hours allowed (SH) to makes 19,300 Jogging Mates? 2. What is the standard labor cost allowed (SH × SR) to make 19,300 Jogging Mates? 3. What is the labor spending variance? 4. What is the labor rate variance and the labor efficiency variance? 5. The budgeted variable manufacturing overhead rate is $4.70 per direct labor-hour. During August, the company incurred $51,940 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month. (For requirements 3 through 5, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations.)
Answer #1

SOLUTION

1. Standard labor hours allowed = Actual number of units * Standard hours per unit

= 19,300 * (27 minutes / 60 minutes)

= 8,685 hours

2. Standard labor cost allowed = Standard labor hours allowed * Standard rate per hour

= 8,685 hours * $5.40

= $46,899

3. Labor spending variance = Standard labor cost - Actual labor cost

= $46,899 - $49,158

= $2,259 U

4. Labor rate variance = (Standard rate - Actual rate) * Actual hours worked

= (5.40 - 5.30) * 9,275

= $927.50 F

Actual rate = Direct labor cost / Direct labor hours

= $49,158 / 9,275 = 5.30

Labor rate variance = (Standard hours - Actual hours) * Standard rate

= (8,685 - 9,275) * $5.40

= $3,186 U

5. Variable overhead rate variance = Actual hours worked * (Standard overhead rate - Actual overhead rate)

= 9,275 * (4.70 - 5.60)

= $8,347.50 U

Actual overhead rate = $51,940 / 9,275 = 5.60

Variable overhead efficiency variance = Standard overhead rate * (Standard hours - Actual hours)

= 4.70 * (8,685 - 9,275)

= 2,773 U